OpenStudy (hannaha):

Medal and fan!!! Brendan got 21 questions correct on his social studies test and missed 4 questions. What % did he have correct? R =

OpenStudy (igreen):

Oh, actually if she got 21 correct, and missed 4, then she had a total of 21 + 4 = 25 questions.

OpenStudy (igreen):

Therefore we can do: \(\dfrac{21}{25} \times 100\)

OpenStudy (igreen):

Just divide 21/25, then multiply that by 100.

OpenStudy (hannaha):

84?

OpenStudy (igreen):

Yep!

OpenStudy (hannaha):

thanks!